OpenStudy (anonymous):

I need some help.. I've already done these problems and i got them back and they need fixed.. will someone please help me

OpenStudy (anonymous):

what are they?

OpenStudy (anonymous):

on these three i need to find the surface area

myininaya (myininaya):

for the first one we have 4 triangles and a square. the area of a triangle is .5(base)(height) and the area of a square is (base)^2

myininaya (myininaya):

so we have 4(.5)(base)(height)+b^2

OpenStudy (anonymous):

well i redid the first one and i got 103.47

myininaya (myininaya):

=2(base)(height)+(base)^2=2(4.2)(10)+(4.2)^2 is this what you did?

OpenStudy (anonymous):

yea

myininaya (myininaya):

I got 101.64

OpenStudy (anonymous):

oh yeah that's what i meant..

myininaya (myininaya):

the second we have 4 triangles

myininaya (myininaya):

so we have 4 times .5(b)(h) since there are all congruent triangles

myininaya (myininaya):

=2bh

myininaya (myininaya):

=2(15)(9)=270

myininaya (myininaya):

it looks like you got a check mark on that one and the answer was 67.5?

OpenStudy (anonymous):

yea.. all the ones i post need fixed

myininaya (myininaya):

k

myininaya (myininaya):

does everything make sense so far?

OpenStudy (anonymous):

yeah

myininaya (myininaya):

ok lets look at the last one

OpenStudy (anonymous):

6 times .5bh

myininaya (myininaya):

yes there are 6 triangles and then there is that hexagon hex means 6 right?

OpenStudy (anonymous):

yea

myininaya (myininaya):

lets look at the triangles first the base 6 and we need to find the height

myininaya (myininaya):

thats if we look at one traingle

myininaya (myininaya):

if we divide a triangle and half we can from a right triangle so we can use the Pythagorean thm the hyp is 12 and if we divide the base in half the base is 3 or our newly formed triangle so we can find the height of the triangle let's call it h. we have 3^2+h^2=12^2 implies h=sqrt(135). so the area of the whole traingle is .5(6)(sqrt(135) since we have 6 trangles we have this area times 6 so this equals 18(sqrt(135)) but now we need to find the area of the hexagon

myininaya (myininaya):

if we draw two vertical lines from one vertex to another and one horizontal line from one vertex to another we would form 2 rectangles and 4 traingles

myininaya (myininaya):

looking at one triangle we see that the hyp is 6 and we know that the legs are equal. so we have b^2+b^2=6^2 imples 2b^2=36 implies b^2=18 implies b=sqrt(18) so the area of the triangle is .5(sqrt18)(sqrt(18)=.5(18)=9 since we have 4 of these triangles then we multiply the 9 by 4 =36 now we need to find the area of the rectangles we know 1 side =6 so we need find the length of the other side that runs perpendicular to it

myininaya (myininaya):

but we found that when we were finding the legs of the triangle above so the side running perpendicular to the 6cm side has length sqrt(18) so the area of the rectangle is (6)sqrt(18) but since we have 2 we multiply by 2 giving us 12sqrt(18) now we add all of our areas together

myininaya (myininaya):

so we have the area of the 6 triangles is 18sqrt(135) and we have the area of the hexagon is 12sqrt(18)+36 the total surface area is 18sqrt(135)+12sqrt(18)+36

OpenStudy (anonymous):

so 974.22

myininaya (myininaya):

i got 296.05

OpenStudy (anonymous):

i did my math wrong..

OpenStudy (anonymous):

but thank u

myininaya (myininaya):

that last problem was really long I hope i'm right lol

OpenStudy (anonymous):

can help me with some more

myininaya (myininaya):

more surface area?

OpenStudy (anonymous):

no volume

myininaya (myininaya):

ok i will look at one more and then i have to leave

OpenStudy (anonymous):

the first three are finding volume

myininaya (myininaya):

ok the first one we need to find the area of the base and multiply by the height of the solid

myininaya (myininaya):

the area of the base is .5(b)(h)=.5(6)(8)=24 the height of the solid is 10 so 24(10)=240

myininaya (myininaya):

the second one you shouldnt had divided by 3 it should had been 2

myininaya (myininaya):

the last one you set it up right but you should had got 3451.04

OpenStudy (anonymous):

okay thank you

myininaya (myininaya):

np i hope i was a help

OpenStudy (anonymous):

ur were